$SPAD/src/input e02aef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
nplus1:=5
a:Matrix SF:=
 [[2.0000 ,0.5000 ,0.2500 ,0.1250 ,0.0625 ]]
xcap:=-1.0
result:=e02aef(nplus1,a,xcap, -1)

———-

References^{[1] nothing}

4